OpenStudy (anonymous):

Part 1: Explain, in complete sentences, the difference between a radical equation and an equation simply containing a radical. Part 2: Provide an example of each.

OpenStudy (anonymous):

In a radical equation, at least one variable of the equation is under a radical sign, or (equivalently) has a fractional exponent. Example: sqrt(x+1) = 25 Or (x+1)^(-1/2) = 25 which is the same equation, just expressed differently An equation with just a radical in it could have the radical (root) of a number or a constant involved in the terms of the equation. y^3 +sqrt(A) = sqrt(7) is an example of such an equation, since the variable y does not have a fractional power (equivalent to having a radical sign over it).

OpenStudy (anonymous):

I found this answer on http://answers.yahoo.com/question/index?qid=20110327182504AA4fLFZ Just so you know:)